OpenStudy (anonymous):

Solve each question using substitution or elimination: x-2y=-18 and 3x+4y=6

OpenStudy (ankit042):

You can start solving this by multiplying eq two by 3 and then subtracting the two

OpenStudy (anonymous):

why?

OpenStudy (ankit042):

We are trying to solve for y first. When you proceed with the method i mentioned above you can convert your equation in two variables to equation in a single variable. Which you can solve easily

OpenStudy (anonymous):

Ok I got a possible answer which is x=5, did I answer the question?

OpenStudy (ankit042):

please recheck your answer

OpenStudy (anonymous):

What did I do wrong

OpenStudy (anonymous):

x=-6?

OpenStudy (anonymous):

y = 6 and x= -6 because: x= -18+2y so 3(-18+2y) +4y=6 -54+6y+4y =6 10y =60 y=6

OpenStudy (anonymous):

ok thanks a bunch

OpenStudy (ankit042):

I think x is -4